OpenStudy (anonymous):

Say that a and b are the lengths of the legs of a right triangle and c is the length of the hypotenuse. Find a if b = 6 and c = 10.

OpenStudy (anonymous):

11.62

jhonyy9 (jhonyy9):

so hope you know pythagora what say that hypothenuse c squared is equale the sum of square a and square b c^2 = a^2 +b^2 100 = a^2 + 36 a^2 = 100 -36 = 64 a =8 hope this is understandably good luck bye